OpenStudy (anonymous):

f(x)=ln[(x^2+1)/(x^2-1)] , find f'(x)

OpenStudy (anonymous):

\[\frac{d}{dx} \ln(x^2+1) - \frac{d}{dx} \ln(x^2-1) \] since ln(a/b) = lna-lnb

OpenStudy (anonymous):

first one =\[\frac{2x}{x^2+1}\] second one =\[\frac{2x}{x^2-1}\]

OpenStudy (anonymous):

so\[f'(x) = \frac{2x}{x^2+1} - \frac{2x}{x^2-1}\]

OpenStudy (anonymous):

Thanks!

OpenStudy (anonymous):

=)

OpenStudy (waheguru):

sarah i need ur help

OpenStudy (waheguru):

and hero